# Questioning the Survey Responses of Large Language Models

### v3

@@ E.2 GAS and WVS surveys
We obtain the ATP survey questions and their corresponding human responses from the GlobalOpinionsQA repository.131313https://huggingface.co/datasets/Anthropic/llm_global_opinions When comparing the similarity of models’ responses to the population-level survey responses of different countries, we use the countries and the similarity metric considered by Durmus et al., [2023]. We find that all models produce survey responses that are more similar to those of the uniformly random baseline than to those of any of the demographic subgroups, see Figure 17.
@@ E.3 Relative alignment for ATP and GAS/WVS surveys
Note that Santurkar et al., [2023] observe that RLHF can result in a “substantial shift […] towards more liberal, educated, and wealthy [demographic groups]”. Our results suggest that this could be an artifact of systematic biases. For the ATP surveys, we observe one outlier for which its alignment before adjustment is not correlated with the entropy of subgroup’s responses: Llama 2 70B Chat, an RLHF-tuned model. However, after adjustment, Llama 2 70B Chat’s alignment trend is remarkably similar to that of Llama 2 70B and all other LLMs, see Figure 19.
@@ E.4 ANES survey
We present questions in the multiple-choice format described in Section 2, using the Interviewer:, Me: prompt style described by Argyle et al., [2022]. We retrieve the 2016 ANES data from the official website141414https://electionstudies.org/data-center/2016-time-series-study/, and process it such that it matches in form the questionnaire designed by Argyle et al., [2022]. We find that the trained classifiers can discriminate between the model-generated data and the ANES data with very high accuracy (99%), see Figure 20.

### v4

@@ E.2 GAS and WVS surveys
We obtain the ATP survey questions and their corresponding human responses from the GlobalOpinionsQA repository.111111https://huggingface.co/datasets/Anthropic/llm_global_opinions When comparing the similarity of models’ responses to the population-level survey responses of different countries, we use the countries and the similarity metric considered by Durmus et al., [2023]. We find that all models produce survey responses that are more similar to those of the uniformly random baseline than to those of any of the demographic subgroups, see Figure 14.
@@ E.4 ANES survey
We present questions in the multiple-choice format described in Section 2, using the Interviewer:, Me: prompt style described by Argyle et al., [2023]. We retrieve the 2016 ANES data from the official website121212https://electionstudies.org/data-center/2016-time-series-study/, and process it such that it matches in form the questionnaire designed by Argyle et al., [2023]. We find that the trained classifiers can discriminate between the model-generated data and the ANES data with very high accuracy (99%), see Figure 15.
Appendix F Sequential sampling of responses
Motivated by recent findings of Argyle et al., [2023] we conducted an additional investigation where we seek to fill entire ACS questionnaires in a sequential manner, in order to generate for each language model a synthetic dataset of responses. This data emulates in form the ACS dataset collected by the U.S. Census Bureau. We then study the extent to which such synthetic datasets resemble the ACS dataset.
F.1 Methodology
We present survey questions in the same order as in the ACS questionnaire. When querying a model to answer survey question , we include a summary of the previously sampled answers in context. 131313The maximum number of tokens in a filled questionnaire was less than 1024 tokens in all cases, thus fitting entirely within the context window of all surveyed models. We then sample from the model’s output probability distribution over answers, and continue to the next question. We illustrate this sequential process in Figure 16. We refer to Appendix D.3 for results collected with different variations of how a model’s previous answers are integrated into the prompt. We find our results to be robust to these prompt variations.
F.2 The discriminator test
F.3 Contrast with silicon samples
Argyle et al., [2023] propose “silicon sampling”, a methodology to produce synthetic survey respondents using LLMs by conditioning on actual survey respondents. They focus on a subset of 12 questions from the 2016 American National Election Studies (ANES) survey. For every human respondent, they construct a corresponding “silicon individual” by querying GPT-3 to predict the ANES respondent’s answer to each survey question given the respondent’s answers to all other questions. Their results indicate that, for the 2016 ANES survey, GPT-3 can be a fairly calibrated predictor of an individual’s answer to some survey question conditioned on the respondent’s answers to all other survey questions.141414Lee et al., [2023] and Sanders et al., [2023] study imputation tasks similar to those of Argyle et al., [2023], and find that LLMs are not calibrated predictors for a variety of such tasks.
However, Argyle et al., [2023] emphasize that important insights can be gained by emulating the survey responses of human populations “prior to or in the absence of human data”. In this work we have considered precisely the setting where models’ responses are obtained in the absence of human data.151515Conceptually, to generate each synthetic individual, we start with a blank survey questionnaire and prompt the LLM to sequentially fill the entire questionnaire. Whereas we only prompt LLMs with their own responses (i.e., to previous survey questions), Argyle et al., [2023] prompt the model only with actual human responses from the 2016 ANES data. To investigate how our findings transfer to the ANES, we reproduce the experiments of Section F using the 2016 ANES survey questionnaire considered by Argyle et al., [2023] and their “interview-style” prompt. We apply the discriminator test, and find that the trained classifiers can discriminate between the model-generated data and the ANES data with accuracy % (see Appendix E), indicating that models’ responses are markedly different to those in the ANES data.